1/5 + 1/11 = please break down how to do this

Answer by stanbon(75887)

1/5 + 1/11
----
least common denominator = 5*11 = 55
-----
Rewrite each fraction with the lcd as its denominator:
----
11/lcd + 5/lcd
----
Combine the nererators over the lcd:
= (11+5)/lcd
-----
= 16/55
====================
Cheers,
Stan H.